OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (10, 12) and is perpendicular to 5x + 8y = 1 Please show all of your work

OpenStudy (nenadmatematika):

y=-5/8x+1/8 is the equation of the given line written in slope-intercept form. Now the slope of the line that is perpendicular to this line has the slope: -1/(-5/8)=8/5 So now you have the slope and the point (10,12) which is sufficient for writing its equation: y-12=8/5(x-10) so y=(8/5)x-4